OpenStudy (anonymous):

Alicia rows 7 miles downstream in the same time she rows 3 miles upstream. She rows downstream 3 miles per hour faster then she rows upstream. Find the rowing rate for each way.

OpenStudy (anonymous):

lets say let t = time taken for her to row in the river x = downstream speed/hr y = upstream speed/hr so we have: distance = time * velocity 7 = t*x 3 = t*y x = 3+y solve now for x, y, t t = 3/y ; y= 3/t t = 7/x ; x = 7/t 7/t = 3 + 3/t 7 = 3t + 3 3t = 4 t = 4/3 y = 3/(4/3) = 9/4 miles/hr x = 7/(4/3) = 21/4 miles/hr

OpenStudy (anonymous):

thankyou!!! so downstream speed is 21 divided by 4 orrrrr am i wrong?